OpenStudy (anonymous):

Please help! How do I work out y = 2x + 5?

OpenStudy (katrinakaif):

You mean graph it?

OpenStudy (anonymous):

Yes. I need to find the slope

OpenStudy (katrinakaif):

Okay Pupply Luv I will help.. First off: 1) In what format is this equation in?

OpenStudy (anonymous):

It says: What is the slope? y = 2x + 5

OpenStudy (katrinakaif):

Okay..So what FORMAT is this equation in? I will guide you through this.. =)

OpenStudy (anonymous):

lol sorry! But what do you mean by the format?

OpenStudy (katrinakaif):

Okay..all of slope equations..are in the format Y = MX + B ... Do you know what "M" is? Is it the slope or the y intercept?

OpenStudy (anonymous):

It's the slope.

OpenStudy (katrinakaif):

Good..Now let us refer to your given equation which is y = 2x + 5 in the format y = mx + b Since we know M is the slope..what is "m" in the equation y = 2x + 5?

OpenStudy (anonymous):

2 is m

OpenStudy (katrinakaif):

You found your slope!! =)

OpenStudy (anonymous):

:O Thanks!

OpenStudy (katrinakaif):

You have done it yourself! =) Congratz ;)